chuck and dana agree to meet in chicago for the weekend. Chuck travels 148 miles in the same time that dana travels 124 mile. If Chuck;s rate of travel is 6mph more than Dana's then at what rate does chuck travels.?
Answer by mananth(12270)

let Dana's rate of travel be xmph
Chuck travels at a rate of x+6 mph
.
Time taken by Dana to travel 124 miles = 124/x
Time taken by chuck in the same time = 148/x+6
124/x = 148/x+6
124(x+6)=148x
124x + 744=148x
148x-124x= 744
24x= 744
x=31 which is dan'a rate of travel
chuck travels 31 +6 = 37 mph